OpenStudy (anonymous):

Use a graph of f or some other method to determine what, if any, value to assign to f(a) to make f continuous at x = a. f(x)= (x) / (7x^2-x) ; a= 0

OpenStudy (anonymous):

Yes a=-1

OpenStudy (anonymous):

Use a graph of f or some other method to determine what, if any, value to assign to f(a) to make f continuous at x = a. f(x)= (x) / (7x^2-x) ; a= 0 Slution: a) substitution with x=a=0 shows non defined fnc. b) factorize the equation further: x/(x(7x -1)) = (x/x)*(1/(7x-1)) =1/(7x-1) Now substitute x=a=0 => 1/(7x-1) = -1. Therefore if we substitute for a=-1 we have a continuos fnc defined at x=-1.

OpenStudy (anonymous):

is that ok?